Given the following JavaScript code, what will be displayed on the web page?

var x = 5;

var y = 12;

while (x < y)

{

     document.write(x);

     document.write("<br>");

     x=x+3;

}
Answer #1

Here is the output produced :

5
8
11


Explanation:

While loop executes until the given condition is true .
ie) in this case until x is less than y.


1 st Iteration :

It prints 5 since 5 is less than 12 .
<br> means break line which is new line.

x=x+3
This means
x = 5+3 = 8 now .
x is 8 now .


2 ns iteration
It prints 8.
new line.

x= 8+3 = 11
x is 11 now.

3rd iteration :

11 is leass than 12 hence .

it prints 11.

x= 11+3

x is 14 now.

4th iteration

Is x still less than y?

is 14 less than 12?

No hence program ends.